# A General Simulation-Based Optimisation Framework for Multipoint Constant-Stress Accelerated Life Tests

### v1

@@ 4.1 Linear Life-Stress Relationship
Figure 1: Sample preliminary ALT data depicting a linear trend. Horizontal (red) dashed line shows the experiment duration. Component lifetimes are shown as (grey) open circles. Censoring occurs at $S=0.1$ and $S=0.2$ ($95\%$ and $5\%$ right-censoring respectively). Estimated lifetime at the design stress ($S=0.05$) is shown as solid (red) point.
@@
We consider obtaining individual optimal test plans for $N$-point designs where $N\in\{2,3,4,5,6\}$, and also consider a separate analysis where the optimiser additionally optimises $N$. For each optimisation, the stress level range is $S=[0.10,0.90]$ with a stress granularity of $10^{-5}$, i.e., $S_{j+1}\geq S_{j}+10^{-5}$. A total of $n=100$ units are used in the optimisation, where the minimum number of test units at a given stress level is $n_{j}=1$. (Note that $n$ is a relatively arbitrary number here as it can always be scaled up to reduce the RMSE.) The optimal test plans pertaining to the linear test data (Figure 1) are listed in Table 1. Given that the test plan data is randomly generated, and that the output of DEoptim is itself a random variable, for each test plan, we report the minimum RMSE obtained by DEoptim, along with the mean RMSE and standard error from 100 simulation replicates.
@@
$N$ Stress Levels Min. Mean Std.
@@
6* 0.19 0.90 0.90 6,653 7,083 148
(86) (05) (09)
@@
Referring to Table 1, the optimal test plan (lowest mean RMSE of 7,068) is a 2-point design: $S_{1}=0.2$ $(n_{1}=82)$, $S_{2}=0.9$ $(n_{2}=18)$. When the number of stresses is not specified ($N=6^{*}$ design), the optimiser converges to 2-point design with similar test unit allocation to that of the $N=2$ design. Indeed, for each $N$ provided in Table 1, the optimal high stress is $S_{N}=0.9$, and the optimal low stress is consistently $S_{1}\approx 0.2$. As per Figure 1, $95\%$ of the lifetimes at the minimum stress ($S=0.1$) are right-censored, with much lower censoring ($5\%$) at $S=0.2$. Thus, for this linear setting, it is best to test more units at a lower stress where there is some (but not much) censoring, and fewer units at a much higher stress where there is more information per unit tested (i.e., at the higher stress, units fail much quicker and are more concentrated around the mean lifetime). Overall, it is clear that there are no statistical differences between the optimal test plans for different $N$ values when the RMSE values are considered in relation to the associated standard errors. In Figure 2 we further illustrate these similarities by visualising each optimal test plan configuration.
@@
Figure 2: Comparison of test plan RMSE for linear AFT model. Solid (black) data points show the stress levels. The size of the points are proportional to the test unit allocation. RMSE values for each test plan are shaded in greyscale (dark-to-light shading corresponds to smallest-to-largest RMSE).
@@
Figure 2 shows how, for each $N$, the optimiser converges to $S_{1}\approx 0.2$, $S_{N}=0.9$, with a higher proportion of test units allocated to the lower stress. For $N\in\{3,4,5,6\}$, we see that the optimiser has allocated a very small number of units to some other stresses. Given the complex and stochastic nature of the optimisation problem, it is highly likely that the optimiser has just not yet fully converged to a 2-point design within the given number of iterations (see note in Section 4). In any case, since these additional stresses are associated with such small numbers of test units ($\leq 3$), and the different plans do not differ from each other statistically, all of this points to a 2-point design being optimal.
@@
Next, we assess the performance of some practical variations of the optimal test plan. This provides practitioners with insights into some alternative test plans which, perhaps due to real experimental constraints, might be more convenient/favourable than the optimal test plan. (It should also be worth noting that such real-world physical constraints could also be embedded directly into our optimiser given its generality.) We consider the test plan: $S_{1}=0.2$, $S_{2}=0.9$, and assess various test unit allocations.
@@
Figure 3: Comparison of the optimal test plan against test plans with alternative test unit allocations. Test plan #5 is the optimal test plan.
@@
A comparison of variations of the optimal test plan is given in Figure 3. Here we compare the optimal test plan ($S_{1}=0.2$ ($n_{1}=82$), $S_{2}=0.9$ ($n_{2}=18$)) against test plans where the test units are rounded to multiples of five. When the test units are within ±$\pm 5$ of the optimal allocation (test plans #4 and #6), the performance does not differ statistically from the optimal test plan (plan #5). The average RMSEs of test plans #3 and #7 are higher than the optimal plan, but the performance could still be considered reasonable in some real-world settings. When the test units deviate from the optimal plan by more than ten units (plans #1, #2, and #8), the performance deteriorates considerably. This analysis shows that we can perturb the test plan slightly (to accommodate practical constraints) without any significant loss of information.
@@
Overall, the optimisation results indicate that for Weibull lifetime data with a linear life-stress relationship, the optimal test plan configuration is to use two stress levels: the highest stress, and a low stress at which almost all failures are observed (i.e., there is very little censoring). In terms of test unit allocation, the optimal strategy is to allocate most units (approximately $80\%-90\%$) to the lower stress. These results are consistent with the literature. (See Meeker and Nelson, (1975), Barton, (1991) and Nelson, (2005), for example.)
@@ 4.2 Quadratic Life-Stress Relationship
Figure 4: Sample preliminary accelerated test data depicting a quadratic trend. Horizontal (red) dashed line shows the experiment duration. Component lifetimes are shown as (grey) open circles. Censoring occurs at $S=0.1$ ($80\%$ of lifetimes are right-censored). Estimated lifetime at the design stress ($S=0.05$) is shown as solid (red) point.
@@
$N$ Stress Levels Min. Mean Std.
@@
6* 0.13 0.52 0.90 0.90 0.90 0.90 8,554 8,940 403
@@
Figure 5: Comparison of test plan RMSE for quadratic AFT model.
Both Table 2 and Figure 5 clearly show that the optimiser converges to a 3-point design: $S_{1}=0.13$, $S_{2}\approx 0.5$, $S_{3}=0.9$. As was the case with the linear test plans, the test unit allocation is inversely related to stress level, i.e., the larger the stress, the fewer units allocated. In general, the test plan configurations are very similar; for each $N$, the stress levels are very consistent with only small variation in test unit allocation.
Figure 6: Comparison of RMSE for selected 3-point quadratic designs.
A comparison of selected 3-point quadratic test plans is given in Figure 6, where we can see that rounding numbers to multiples of five, i.e., plans #4 and #6 exhibit no statistical difference from the optimal plan #5.
Overall, in relation to a quadratic Weibull AFT model, the optimal test plan configuration is to use three stress levels: a low stress, an intermediate stress, and a high stress. This coincides with previous works in the literature, for example: Khamis and Higgins, (1996) and Yang, (1994). In relation to test units, the optimal strategy is to allocate more test units to lower stresses; again aligning with the literature (see Nelson, (2005).) The test unit allocation shown in Table 2 closely matches the common rule of thumb to allocate units to the high, medium and low stresses in the ratio 4:2:1 (Zhu and Elsayed,, 2013). More specifically, the optimiser converges to a test unit allocation of $n_{1}=72,n_{2}=22,n_{3}=6$, but were we to apply the conventional 4:2:1 rule, the test unit allocation would be: $n_{1}=57,n_{2}=29,n_{3}=14$. The latter is similar to #2 in Figure 6, which is statistically worse than the optimal. Thus, while our results are in line with broad guidelines provided in the literature, our framework finds a better solution for the specific quadratic relationship considered here.
4.3 Power Law Life-Stress Relationship
Next, we discuss a power law life-stress relationship. Figure 7 depicts preliminary data with a power law trend, so we fit a model of the form: $\mu(S)=\beta_{0}+\beta_{1}\log(S)$, giving estimated coefficients of $\hat{\beta_{0}}=-6.9,\hat{\beta_{1}}=-6.2,\hat{\sigma}=0.5$.
Figure 7: Sample preliminary accelerated test data depicting a power law trend. Horizontal (red) dashed line shows the experiment duration. Component lifetimes are shown as (grey) open circles. All lifetimes are uncensored. Estimated lifetime at the design stress ($S=0.05$) is shown as solid (red) point.
Table 3: Optimal test plan configurations for power-law AFT model.
$N$ Stress Levels Min. Mean Std.
@@ 4.3 Power Law Life-Stress Relationship
Figure 8: Comparison of test plan RMSE for power law AFT model.
The results are very similar to the previous linear model analysis (see Section 4.1), i.e., the optimiser converges to a 2-point design, where the majority of the test units are allocated to the lower stress. (Meeker and Nelson, (1975) discuss a Weibull power law example with similar results.) The similarity with the linear model is to be expected since the power law is also a two-parameter regression model that is linear on the log-$S$ scale as per Appendix A. The main difference between the linear and power law analyses is the level of the optimal low stress. We saw from the linear model analysis, that the optimal low stress was $S_{1}=0.2$. As per Figure 1, $95\%$ of the lifetimes at $S=0.1$ were right-censored, so the optimiser favoured a higher stress of $S=0.2$. However, with the power law lifetime data (Figure 7), all of the lifetimes at $S=0.1$ are uncensored, and the optimiser has selected this lowest stress as it is not impaired by censoring, while providing more information than $S=0.2$ (in that it is further away from the high stress). This shows that, the proportion of censored lifetimes (dictated by the experiment duration), influences the optimal low stress level of the test plan.
Our analysis of Weibull AFT models for linear, quadratic and power-law functional forms has shown that our proposed simulation-based optimisation framework is sufficiently general to optimise a wide range of test plans. Moreover, where previous literature has provided analytic results/guidelines, our optimiser provides similar results, while being more targeted to the specific parameter values at hand. As discussed previously, the main advantages of our approach is that it offers: (i) capability to handle more complicated test plans where analytical methods are not feasible, and (ii) flexibility to be applied across a range of practical ALT experiments. To this end, in the next section we utilise a case study to demonstrate how the framework is easily applied to real-world accelerated test data.
5 Case Study - Electronic Component Lifetime
We now demonstrate the practical application of our optimisation framework by utilising industrial ALT data from an industry partner of ours. In this study, the manufacturing company are interested in estimating the lifetimes of electronic components whose intended operating voltage is 2.1kV (kilo volts). An initial accelerated life test was already conducted where $n=23$ components were tested at the each of the following (constant) voltage stresses: $\{2.5,2.75,3,3.5,4,4.5,5\}$kV. All components were tested for a period of 6 months, or until failure, whichever came first. The lifetime data from this experiment, and a selection of fitted Weibull reliability models (linear, quadratic and power law), are given in Figure 9.
Figure 9: Constant-stress accelerated test data with fitted Weibull reliability models. Horizontal (red) dashed line shows the experiment duration. Component lifetimes are shown as (grey) open circles. All lifetimes at 2.5kV and 2.75kV are right censored. Estimated design-stress ($2.1$kV) lifetimes for each model are shown as a solid (red) point.
Of the three Weibull models fitted, the quadratic model is the best fit (based on the Akaike information criterion), so we use this model in our analysis. The estimated model parameters are: $\hat{\beta_{0}}=30.310,\hat{\beta_{1}}=-10.108,\hat{\beta_{2}}=0.858,\hat{\sigma}=0.2418$. These parameters were then used within our optimiser.
$N$ Voltage Levels Min. Mean. Std.
$V_{1}$ $V_{2}$ $V_{3}$ $V_{4}$ $V_{5}$ $V_{6}$ RMSE RMSE Error
3 2.91 3.96 5.00 34,174 37,292 984
(75) (68) (18)
4 2.94 3.96 5.00 5.00 35,910 39,090 1,334
(78) (60) (14) (09)
5 2.89 3.93 4.06 5.00 5.00 36,112 37,397 931
(85) (61) (03) (02) (10)
6 2.91 4.12 5.00 5.00 5.00 5.00 36,449 37, 892 920
(72) (65) (05) (05) (12) (02)
6* 2.88 4.00 5.00 5.00 5.00 5.00 34,379 37,259 942
(79) (60) (01) (05) (06) (10)
Figure 10: Comparison of test plan RMSE for quadratic AFT model (industrial data).
The optimisation results (Table 4 and Figure 10) show that the optimal test plan is a 3-point plan: $V_{1}\approx 2.9$kV, $V_{2}=4$kV and $V_{3}=5$kV. From a total of $n=161$ components, approximately 80 are tested at 2.9kV, 60 at 4kV, and the remaining 21 at 5kV. We can see from Figure 11 that, in general, modest variations (5 units or so) from the optimal component allocation (test plan #4) provide comparable performance.
@@ 5 Case Study - Electronic Component Lifetime
Overall, the findings from this application align broadly with our previous (simulated) quadratic model analysis (Section 4.2) and the 4-2-1 rule from previous literature. Our Section 4.2 results assign approximately 70%, 20%, and 10% to each of a low, medium, and high stress, whereas the 4-2-1 rule implies allocations of 57%, 29%, and 14% respectively. For the electronic component data in the current section, the allocation is 50%, 37%, and 14%, respectively, to voltages $V_{1}=2.9$kV (low), $V_{2}=4$kV (medium), and $V_{3}=5$kV (high). Therefore, from a very generic perspective, the guideline is that more devices are allocated to lower stresses (as was also the case for the linear and power law relationships). While this general advice may be sufficient in some situations, the specific proportions and positions of the stresses will depend on the exact problem at hand. Our optimisation framework can be applied to configure any application-specific test plan, and such a plan can indeed be statistically superior than using only broad guidelines.

### v2

@@ 4.1 Linear Life-Stress Relationship
Figure 1: Sample preliminary ALT data depicting a linear trend. Horizontal (red) dashed line shows the experiment duration, $t_{e}$. Component lifetimes are shown as (grey) open circles. Censoring occurs at $S=0.1$ and $S=0.2$ ($95\%$ and $5\%$ right-censoring respectively). Estimated lifetime at the design stress ($S=0.05$) is shown as solid (red) point.
@@
We consider obtaining individual optimal test plans for $N$-point designs where $N\in\{2,3,4,5,6\}$, and also consider a separate analysis where the optimiser additionally optimises $N$. For each optimisation, the stress level range is $\bm{S}=[0.10,0.90]$ with a minimum stress distance (granularity) $D$, such that $S_{j+1}\geq S_{j}+D$. In this work, we specify a small value of $D=10^{-5}$, which, from a practical perspective, essentially allows a given $N$-point design the ability to reduce to a lower-than-$N$-point design. For example, for the $N=6$ design shown in Table 1, stresses $S_{3},S_{4},S_{5}$ and $S_{6}$ are each separated by a distance of $D=10^{-5}$ which, from a practical perspective (rounding to two decimal places), can be considered as a single stress $S_{3}=0.90$. In some situations this property may not be desirable, for example, a practitioner might need to more strictly enforce an $N$-point design, in which case $D$ can set to be a larger number, e.g., to align with the stress resolution of their physical experimental setup. In Appendix B.1, we provide optimisation results with the constraint that $D=10^{-1}$, which then creates a bigger separation between neighbouring stress values.
In terms of experimental test units, we consider a total of $n=100$ in the optimisation, where the minimum number of test units at a given stress level is $n_{j}=1$. Note that $n$ is a relatively arbitrary number here as it can always be scaled up to reduce the RMSE. Indeed, for the linear AFT scenario, the RMSE is reduced by approximately $30\%$ when the number of test units is increased from $n=100$ to $n=200$; this is shown in Appendix B.2.
We remind the reader of the stochastic nature of the overall optimisation framework: DEoptim is a stochastic optimiser, which is minimising a stochastic objective function (i.e., Monte-Carlo-calculated RMSE). Therefore, the test plan results will naturally vary (slightly) from one run of the algorithm to another. To account for this variability, in addition to reporting the minimum RMSE obtained by DEoptim, we also report the mean, standard error, and the 5th and 95th percentiles of the RMSE following 1000 simulation replicates at the given DEoptim test plan. These additional metrics are important for understanding whether or not the given plans differ statistically from each other. (Of course, one can also run DEoptim itself multiple times, and we have run some variations of the linear-AFT setup provided in Appendix B, all of which display similar results to Table 1 in terms of best test plan.)
@@
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
@@
6* 0.19 0.27 0.90 0.90 0.90 0.90 6,455 7,070 166 6,788 7,353
(82) (01) (01) (01) (01) (14)
The optimal test plans pertaining to the linear test data (Figure 1) are listed in Table 1, indicating that a 2-point design is best. In particular, the 2-point design has the lowest mean RMSE (7,038), and suggests $n_{1}=85$ units at stress $S_{1}=0.19$ and $n_{2}=15$ units at stress $S_{2}=0.89$. Given the values of standard errors and percentiles, it is clear that this design does not differ statistically from the other $N$-point designs. Moreover, given the small $D=10^{-5}$ value we have used, from a practical perspective, all designs essentially collapse to a 2-point design where most test units are assigned to the low stress (of around 0.2) with the remaining units assigned to the high stress (of around 0.9).
@@
As per Figure 1, $95\%$ of the lifetimes at the minimum stress ($S=0.1$) are right-censored, with much lower censoring ($5\%$) occurring at $S=0.2$. Thus, for this linear setting, it is best to test more units at a lower stress where there is some (but not much) censoring, and fewer units at a much higher stress where there is more information per unit tested (i.e., at the higher stress, units fail much quicker and are more concentrated around the mean lifetime).
@@
Figure 2: Comparison of optimal test plans and RMSE for linear AFT model. Solid (black) data points show the stress levels. The size of the points are proportional to the test unit allocation. RMSE values for each test plan are compared using greyscale shading, where light-to-dark shading (white-to-black) corresponds to the smallest mean RMSE plus/minus three times the associated standard error. (But all designs end up with the same shade of grey as they differ so little statistically.)
@@
Figure 2 provides a visualisation of the results of Table 1, where for each $N$, the optimiser converges to $S_{1}\approx 0.2$, $S_{N}\approx 0.9$, with a higher proportion of test units allocated to the lower stress. For $N\in\{5,6,6^{*}\}$, we note that the optimiser has allocated a very small number of units to other stresses. Given the complex and stochastic nature of the optimisation problem, it is likely that the optimiser has not yet fully converged to a 2-point design within the given number of iterations. From a practical perspective of physical testing, it is unlikely that one would run a given design point with such a small number of units, when not supported by any meaningful improvement in RMSE. More precisely, reallocating these “spurious” few units to the smaller and larger stresses collapses to the 2-point design, which yields the same RMSE from a statistical perspective (as indicated by the grey shading in the plot).
@@
Next, we assess the performance of some practical variations of the optimal test plan. This provides practitioners with insights into some alternative test plans which, perhaps due to real experimental constraints, might be more convenient to implement. (It is worth noting that such real-world physical constraints can also be embedded directly into our optimiser as the data can be simulated subject to any constraints.)
@@
Figure 3: Comparison of 2-point linear test plans with various test unit allocations and/or rounding of stresses. Test plan # 1 is the optimal test plan.
@@
A comparison of variations of the optimal test plan is given in Figure 3. Here we compare the optimal test plan ($S_{1}=0.19$, $n_{1}=85$; $S_{2}=0.89$, $n_{2}=15$) against several test plans where the test units are rounded (plans #2, #3 and #4), and stresses are also rounded (plans #5, #6 and #7). Using the optimal stresses, when the test units are within $\pm 5$ of the optimal allocation (test plans #2 and #3), the performance does not differ much from the optimal test plan (plan #1). However, when the test units deviate by $\pm 10$ from the optimal, the RMSE does increase significantly. Referring to the test plans where both test units and stresses are rounded, the average RMSE for test plans #5 and #6 are very similar to the optimal plan. However, as with test plan #4, the RMSE of test plan #7 is larger. This analysis shows that we can perturb the test plan slightly (to accommodate practical constraints) without any significant loss of performance.
@@
Overall, the optimisation results indicate that for Weibull lifetime data with a linear life-stress relationship, the optimal test plan configuration is to use two stress levels: the highest stress, and a low stress at which almost all failures are observed (i.e., there is very little censoring). In terms of test unit allocation, the optimal strategy is to allocate most units (approximately $85\%$) to the lower stress. These results are consistent with the literature. (See Meeker:1975, Barton:1991 and Nelson:2005, for example.)
@@ 4.2 Quadratic Life-Stress Relationship
Figure 4: Sample preliminary accelerated test data depicting a quadratic trend. Censoring occurs at $S=0.1$ ($80\%$ of lifetimes are right-censored).
@@
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
@@
6* 0.13 0.53 0.58 0.90 0.90 8,239 9,062 219 8,717 9,429
(75) (16) (01) (01) (07)
Table 2 shows that the optimiser converges to a 3-point design: $S_{1}=0.13$, $S_{2}\approx 0.5$, $S_{3}=0.9$. Akin to the linear test plans, the test unit allocation is inversely related to stress level, i.e., the larger the stress, the fewer units allocated. We can see that the test plans are very similar: for each $N$, the stress levels are reasonably consistent with relatively small variation in test unit allocation. As shown in Figure 5, each test plan exhibits a similar grey shade indicating that there is little statistical difference between test plans. (Recall that the greyscale shading is based on the standard error of the RMSE.) For $N=\{5,6,6\textsuperscript{*}\}$, the optimiser identifies a fourth stress with a small number of test units ($n\leq 2$). Given the small number of test units, it seems reasonable (from the point of view of physical testing), to consider these as 3-point plans. Overall, given that the test plans are not statistically different, the optimisation results indicate that a test plan with three stresses is optimal.
Figure 5: Comparison of optimal test plans and RMSE for quadratic AFT model; greyscale colouring as per Figure 2.
A comparison of 3-point quadratic test plans is given in Figure 6. Test plans #2, #3 and #4 have rounded test units, and test plans #5, #6 and #7 have both test units and stresses rounded. We can see that slight rounding of test units (plans #2 and #3) provides comparable performance to the optimal plan (plan #1), whereas the other deviations presented here (in plans #4, #5, #6, #7) lead to larger increases in RMSE.
Figure 6: Comparison of 3-point quadratic test plans. (Test plan # 1 is the optimal plan.)
Overall, in relation to a quadratic Weibull AFT model, the optimal test plan configuration is to use three stress levels: a low stress, an intermediate stress, and a high stress. This coincides with previous works in the literature, for example: Khamis:1996 and Yang:1994. In relation to test units, the optimal strategy is to allocate more test units to lower stresses; again aligning with the literature (see Nelson:2005.) The test unit allocation shown in Table 2 is in line with the common rule of thumb to allocate units to the low, medium and high stresses in the ratio 4:2:1 (Zhu:2013). More specifically, the optimiser converges to a test unit allocation of $n_{1}=71,n_{2}=24,n_{3}=5$, but were we to apply the conventional 4:2:1 rule, the test unit allocation would be: $n_{1}=57,n_{2}=29,n_{3}=14$. Whilst our results are reasonably consistent with broad guidelines provided in the literature, our framework finds a solution tailored to the specific quadratic relationship being analysed.
4.3 Power Law Life-Stress Relationship
Next, we discuss a power law life-stress relationship. Figure 7 depicts preliminary data with a power law trend, so we fit a model of the form: $\mu(S)=\beta_{0}+\beta_{1}\log(S)$, giving estimated coefficients of $\hat{\beta_{0}}=-6.9,\hat{\beta_{1}}=-6.2$ and $\hat{\sigma}=0.5$.
Figure 7: Sample preliminary accelerated test data depicting a power law trend. All lifetimes are uncensored.
Table 3: Optimal test plan configurations for power law AFT model.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.10 0.89 7,239 7,945 191 7,641 8,274
(77) (23)
3 0.10 0.56 0.90 7,331 7,953 182 7,656 8,251
(77) (01) (22)
4 0.10 0.81 0.90 0.90 7,303 7,902 184 7,578 8,198
(81) (03) (15) (01)
5 0.10 0.86 0.90 0.90 0.90 7,388 7,895 174 7,619 8,183
(77) (01) (19) (01) (02)
6 0.10 0.70 0.90 0.90 0.90 0.90 7,322 8,006 186 7,699 8,311
@@ 4.3 Power Law Life-Stress Relationship
Figure 8: Comparison of optimal test plans and RMSE for power law AFT model; greyscale colouring as per Figure 2.
As shown in Table 3 and Figure 8, the results from the power law model analysis are very similar to the previous linear model analysis (see Section 4.1), i.e., the optimiser converges to a 2-point design, where the majority of the test units are allocated to the lower stress. (Meeker:1975 discuss a Weibull power law example with similar results.) The similarity with the linear model is to be expected since the power law is also a two-parameter regression model that is linear on the log-$S$ scale as per Appendix A. The main difference between the linear and power law analyses is the level of the optimal low stress. We saw from the linear model analysis, that the optimal low stress was $S_{1}=0.19$. As per Figure 1, $95\%$ of the lifetimes at $S=0.1$ were right-censored, so the optimiser favoured a higher stress of $S=0.19$. However, with the power law lifetime data (Figure 7), all of the lifetimes are uncensored, and the optimiser has selected the lowest stress of $S=0.1$ as it is not impaired by censoring, while providing more information than $S=0.19$ (in that it is further away from the high stress). This shows that the proportion of censored lifetimes (dictated by the experimental duration) influences the optimal low stress level of the test plan. Appendices B.4 and B.5 show that, for the linear AFT model, an optimal low stress of $S_{1}=0.23$ is obtained when the experiment duration, $t_{e}$, is reduced from one year to 6 months.
Our analysis of Weibull AFT models for linear, quadratic and power law functional forms has shown that our proposed simulation-based optimisation framework is sufficiently general to optimise a wide range of test plans. Moreover, where previous literature has provided analytic results/guidelines, our optimiser provides similar results, while being more targeted to the specific parameter values at hand. As discussed previously, the main advantages of our approach is that it offers: (i) capability to handle more complicated test plans where analytical methods are not feasible, and (ii) flexibility to be applied across a range of practical ALT experiments. To this end, in the next section we utilise a case study to demonstrate how the framework is easily applied to real-world accelerated test data.
5 Case Study - Electronic Component Lifetime
We now demonstrate the practical application of our optimisation framework by utilising industrial ALT data from an industry partner of ours. In this study, the manufacturing company are interested in estimating the lifetimes of electronic components whose intended operating voltage is 2.1kV (kilovolt). An initial accelerated life test was previously conducted where $n=23$ components were tested at each of the following (constant) voltage stresses: $\{2.5,2.75,3,3.5,4,4.5,5\}$kV. All components were tested for a period of 6 months, or until failure, whichever came first. The lifetime data from this experiment, and a selection of fitted Weibull reliability models are given in Figure 9.
Figure 9: Constant-stress accelerated test data with fitted Weibull reliability models. All lifetimes at 2.5kV and 2.75kV are right censored. Estimated design-stress (2.12.1kV) lifetimes for each model are shown as a solid (red) point.
Figure 9 shows that the linear, quadratic and power law models yield significantly different estimates at the design stress of 2.1kV (approximately 5 years, 44 years and 276 years respectively). The contrast in design-stress predictions highlights the importance of model selection within the simulation/optimisation framework. Of course, model selection is a fundamental aspect of many statistical analyses, but it is particularly important in the context of ALT, where model extrapolation is necessary. We can see that the quadratic and power law models appear to give reasonably similar estimates (albeit on the log-scale) at the accelerated stresses, but quite different estimates are obtained once the models are extrapolated to the design stress.
To mitigate such estimation uncertainty, we have used a weighted Akaike information criterion (AIC) approach in this case study (Burnham:2002, Chapter 2). We found that almost all ($99.97\%$) of the AIC weight was allocated to the quadratic model (indicating that the quadratic model provides a significantly better fit compared to the other models). Thus, for this particular analysis, we fit a quadratic model, and use the estimated model parameters: $\hat{\beta_{0}}=30.3,\hat{\beta_{1}}=-10.1,\hat{\beta_{2}}=0.9,\hat{\sigma}=0.2$ within our optimiser. In general, AIC weighted model-averaging is an effective way to account for model selection uncertainty, which could even be propagated into the optimisation algorithm, but that is beyond the scope of this paper.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
@@ 5 Case Study - Electronic Component Lifetime
Figure 10: Comparison of optimal test plans and RMSE for quadratic AFT model (industrial data); greyscale colouring as per Figure 2.
The optimisation results (Table 4 and Figure 10) show that the optimal test plan is a 3-point plan: $V_{1}\approx 2.9$kV, $V_{2}\approx 4$kV and $V_{3}=5$kV, with a corresponding test unit allocation of approximately 80:60:20. We can see from Figure 11 that, variations of 5 test units or so (test plans #2 and #3) and test unit variation coupled with modest stress rounding (test plans #5 and #6), provide comparable performance to the optimal test plan (test plan #1). When the test units allocation is varied further (test plans #4 and #7), the out-of-sample prediction performance is reduced, i.e., the average RMSE is larger. However, in some applications, it could be that these higher-RMSE plans could still be viable, especially if the cost of running the experiment at the largest stress ($S_{3}$) is high. Indeed, the costs associated with the number of units, sizes of stresses, and experimental duration could be all be incorporated into a more general objective function than the purely predictive-performance-based RMSE objective function in (2).
@@
Overall, the findings from this application align broadly with our previous (simulated) quadratic model analysis (Section 4.2) and the 4:2:1 rule from previous literature. Our Section 4.2 results assign 71%, 24%, and 5% to each of a low, medium, and high stress, whereas the 4:2:1 rule implies allocations of 57%, 29%, and 14% respectively. For the electronic component data in the current section, the allocation is 49%, 39%, and 12%, respectively, to voltages $V_{1}\approx 2.9$kV (low), $V_{2}\approx 4$kV (medium), and $V_{3}=5$kV (high). Therefore, from a very generic perspective, the guideline is that more devices are allocated to lower stresses (as was also the case for the linear and power law relationships). While this general advice may be sufficient in some situations, the specific proportions and positions of the stresses will depend on the exact problem at hand. Our optimisation framework can be applied to configure any application-specific test plan, and such a plan can indeed be statistically superior than using only broad guidelines.